two circular cylinders of eaquel volumes have their height in the ratio 9:4 find the ratio of their radii
Answer by mananth(16946)

Let Volume of cylinder be V1
second cylinder be V2
V= pi * r^2*h
V1= pi*r1^2*h1
V2= pi*r2^2*h2
The volumes are equal
So V1=V2
pi*r1^2*h1=pi*r2^2*h2
r1^2/r2^2= h2/h1
r1^2/r2^2= 9/4
r1/r2= 3/2